OpenStudy (sierra_tibbetts):

If g(x) = 2(x − 4), find the value of x if g(x) = 20. 14 32 12 10

OpenStudy (studygurl14):

replace g(x) with 20 and solve for x 20 = 2(x - 4)

OpenStudy (sierra_tibbetts):

So then whats the answer

OpenStudy (mhchen):

The answer is simplifying StudyGurl's equation of 20 = 2(x-4) 10 = x - 4 etc..